OpenStudy (anonymous):

1. Solve √3(x - 2) + 7 = 28 2.What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4?

OpenStudy (bahrom7893):

is the whole thing under sqrt, or is it just the 3?

OpenStudy (anonymous):

(x - 2) is also under it but not the rest

OpenStudy (bahrom7893):

okay so first, simplify a little: sqrt(3(x+2)) = 28-7 = 21 sqrt(3(x+2)) = 21 square both sides: 3(x+2) = 21^2

OpenStudy (bahrom7893):

can u take it up from here?

OpenStudy (anonymous):

yes thank you what about the second question..

OpenStudy (bahrom7893):

Perpendicular means the slope is -1/original slope

OpenStudy (bahrom7893):

So your new slope = -1/(-2) = 1/2

OpenStudy (bahrom7893):

The equation is of the form y = mx + b Plug in the values that we found, and the ones they gave us: 1 = (1/2) * 4 + b solve for b, and you'll end up with the answer.

OpenStudy (bahrom7893):

Well the equation will be: y = (1/2)x + b, whatever b is